### v1

@@ 1 Introduction
The challenge of attributing outcomes to early stages under terminal-only observation has gained particular urgency in AI, where large language models generate extended reasoning chains but only the final answer can be verified. Empirically, value networks trained on outcome rewards “collapse”: they converge to near-constant predictions for early steps, as if outcomes carry no information about the chain’s beginning (Kazemnejad et al. 2024). Process reward models that provide step-level supervision dramatically improve performance (Lightman et al. 2024), but require expensive human annotation. When is such annotation necessary, and when does outcome supervision suffice?
This paper answers these questions by establishing information-theoretic foundations for credit assignment. Using Strong Data Processing Inequalities, we prove that signals connecting early steps to terminal outcomes decay exponentially with depth. The analysis reveals a phase transition: there exists a critical horizon $H_{\mathrm{crit}}$ such that steps within this distance of the outcome can be evaluated from endpoint data, while steps beyond cannot, regardless of sample size, algorithm, or computational budget. This is not an algorithmic limitation but an information-theoretic impossibility. The critical horizon scales logarithmically with sample size and signal strength but inversely with a per-stage contraction rate that captures how much each processing step blurs the distinction between good and bad earlier steps. In AI reasoning systems, sampling temperature controls this blurring: higher temperature produces more diverse continuations, making it harder to trace outcomes back to early steps. In manufacturing, the analogous driver is the overlap among downstream quality-grade distributions: variability-reducing steps (polishing, annealing, rework), pooling operations (lot mixing, kitting, assembly), and coarse metrology bins all cause products of different upstream grades to produce similar distributions of grades at the next stage, fading the signal about the original grade. The critical horizon formula thus determines whether intermediate feedback is a cost-saving measure or a feasibility requirement.
@@
This many-to-one mapping from observations to abstract states is the mechanism behind the per-stage contraction introduced above: when multiple observable states collapse to the same abstract state, transitions from distinct starting points lead to overlapping distributions over next states, and statistical distinguishability shrinks. This occurs even when raw dynamics are deterministic (as with token generation in language models), because the induced dynamics on the abstract space contract. The implication for learning systems is immediate: when signal decay is severe, outcome distributions conditioned on different early states become nearly identical, and training data contains almost no information about which early states are better. The critic collapse phenomenon observed in AI systems (Kazemnejad et al. 2024) is thus statistically optimal behavior given the available information, not a bug to be fixed but a symptom of a fundamental limit.
A natural response to signal decay is to generate more observations: multiple parallel trajectories from the same starting point, as in replicated testing or width-based methods like GRPO (DeepSeek-AI 2024). We show this strategy provides only limited relief: correlation among rollouts caps the effective width at $1/\rho$, where $\rho$ is the pairwise correlation between rollout outcomes, and adding more rollouts yields diminishing returns.
@@
We develop a detailed application to large language model reasoning (Section 14 in the Electronic Companion), where the state abstraction maps token sequences to semantic content and sampling temperature controls the contraction coefficient. We validate the theoretical predictions through numerical experiments on both synthetic Markov chains and LLM reasoning chains on GSM8K (a grade-school math benchmark).
@@
The framework yields concrete guidance for both AI and operations practitioners. For AI systems, it explains why process reward models improve performance (they provide signal beyond the critical horizon), why methods like VinePPO (Kazemnejad et al. 2024) succeed (they reduce effective depth below $H_{\mathrm{crit}}$), and why width-based methods face diminishing returns (correlation caps effective width). For operations managers, it provides a direct diagnostic. Consider a 50-stage semiconductor process. With contraction rate $\eta=0.85$, signal strength $\Delta^{2}=0.2$, and 10,000 test wafers, $H_{\mathrm{crit}}\approx 47$: at least one intermediate checkpoint is required. With $\eta=0.90$, the critical horizon extends to 72 steps, and terminal testing alone suffices. (In practice, $\eta$ is auditable: estimate adjacent-stage grade transition matrices from production data and compute the overlap-based bounds in Section 3.2.)
@@ 1.1 Literature Review
The information-theoretic structure underlying credit assignment connects five research streams that have developed largely independently: inspection design in operations research, credit assignment in AI reasoning, reinforcement learning theory, information-theoretic analysis of sequential processes, and the emerging application of OR tools to AI system design.
Inspection design in operations research.
The optimal placement of inspection points in sequential production is a foundational topic in operations research. Lindsay and Bishop (1964) formulated inspection allocation as a serial dynamic program, minimizing total cost of inspection plus penalties for escaping defects. Raz (1986) surveyed extensions incorporating multiple inspection stations and variable costs; Tapiero (1987) introduced imperfect inspection and learning effects. Recent work incorporates adaptive inspection (Tan et al. 2024), predictive maintenance (Compare et al. 2020), and machine-learning-based prognosis (Diez-Olivan et al. 2019). Ghuge et al. (2025) study sequential testing of system components under budget constraints, providing the first constant-factor approximation for nonadaptive score classification. Duenyas et al. (2024) analyze supply chains where components cannot be individually tested, developing coordination mechanisms when stage-level inspection is infeasible, precisely the terminal-only observation regime our theory addresses.
This literature optimizes checkpoint placement to minimize cost under the assumption that defect signals persist through the production chain. Our work identifies a structural limit: in deep processes, signal decay makes observability, not cost, the binding constraint. The critical horizon formula quantifies when this regime applies.
Credit assignment in AI reasoning.
The credit assignment problem was recognized by Minsky (1961) as fundamental to learning systems, yet has resisted systematic theoretical treatment. Our framework provides an information-theoretic characterization of when outcome supervision suffices versus when process supervision becomes necessary, grounded in the signal decay structure of deep sequential processes.
Recent work explores process supervision as a remedy: Lightman et al. (2024) show that step-level labels substantially improve mathematical reasoning; Uesato et al. (2022) compare outcome and process supervision for mathematical reasoning; Wang et al. (2024) develop automated step-labeling methods. These studies demonstrate that process supervision helps but do not characterize when it becomes necessary. Our Signal Decay Bound fills this gap.
Empirically, value networks trained on outcome data exhibit “critic collapse”, converging to near-constant predictions on long chains (Kazemnejad et al. 2024). Width-based methods such as GRPO (DeepSeek-AI 2024) address this by replacing learned critics with Monte Carlo estimation. Our framework explains both phenomena: critic collapse reflects the absence of signal beyond the critical horizon, and width-based methods face diminishing returns because correlation caps effective width.
Reinforcement learning theory.
Standard reinforcement learning theory establishes polynomial sample complexity under regularity conditions such as bounded concentrability (Kakade 2003, Azar et al. 2013). Recent work by Jia et al. (2025) shows that outcome and process supervision are polynomially equivalent when concentrability is bounded. Our setting falls outside this regime: in deep sequential processes where each step may fail, concentrability grows exponentially in horizon, and the polynomial equivalence breaks down. This is precisely why the critical horizon matters. Sample complexity in related RL settings has received attention in the OR community: Hu et al. (2025) establish fast convergence rates for offline RL, Cheung et al. (2023) develop algorithms for nonstationary environments, and Lykouris et al. (2025) address exploration under adversarial corruption. These works assume bounded concentrability or benign nonstationarity; the deep sequential regime we study, where signal decays exponentially, falls outside their scope.
A second point of departure from standard reinforcement learning concerns objective structure. Standard algorithms aggregate rewards additively, but sequential validity requires all steps to be correct, inducing a multiplicative objective. This connects to risk-sensitive MDPs (Howard and Matheson 1972), multiplicative-reward models (Baier et al. 2025), and minimax-optimal risk-sensitive RL (Wang et al. 2023). We propose a curriculum strategy that interpolates between the learnable additive objective and the correct multiplicative one.
Information-theoretic foundations.
Filter stability results (Del Moral and Guionnet 2001) establish that dependence on initial conditions decays asymptotically in Markov chains, but do not yield the non-asymptotic, closed-form rates that credit assignment requires: a practitioner needs to know how many samples suffice for a process of a given depth, not merely that the problem becomes harder in the limit. Our analysis provides such rates by combining two classical tools. Strong Data Processing Inequalities (Polyanskiy and Wu 2017, Raginsky 2016) quantify per-step information loss, establishing that $\chi^{2}$-divergence contracts by a factor $\eta<1$ at each transition. Le Cam’s method (Le Cam 1973) then translates the resulting divergence bounds into sample complexity lower bounds, yielding the critical horizon formula. Recent extensions to bandits under communication constraints (Pensia et al. 2024) and interactive learning (Chen et al. 2024) address related but distinct settings.
Operations research tools for AI system design.
A growing body of work applies OR methodology directly to AI systems. Snell et al. (2025) frame inference-time compute allocation as an optimization problem, adaptively distributing verification effort across prompts of varying difficulty, an allocation problem that parallels inspection placement at the problem-structure level, though without the information-theoretic signal decay analysis we develop here. Besta et al. (2025) formulate LLM reasoning explicitly within the MDP framework, treating chain-of-thought generation as sequential decision-making. Zhang et al. (2024) apply Monte Carlo Tree Search to guide multi-step reasoning with step-level evaluation, and Tang et al. (2024) model iterative code repair as a multi-armed bandit. Pieroth et al. (2025) use deep RL to compute equilibrium strategies in multistage auctions, and Elmachtoub and Grigas (2022) establish the “predict, then optimize” framework integrating prediction with downstream decision-making, both demonstrating OR-native methodology enriching AI system design. Recent surveys of INFORMS Fellows identify bringing OR rigor to AI challenges as a priority research frontier (Wiberg et al. 2025). Our work contributes to this agenda by showing that Strong Data Processing Inequalities and minimax inspection design, classical OR tools, yield precise, actionable answers to questions the AI community has addressed primarily through empirical heuristics.
@@ 1.2 Paper Organization
Section 6 synthesizes these results into optimal inspection schedules: uniform spacing under homogeneous contraction and a greedy information-distance algorithm under heterogeneous contraction. Section 7 validates the theoretical predictions through synthetic Markov chains and LLM reasoning chains on GSM8K, covering all four main results. The Electronic Companion develops a detailed application to language model reasoning (Section 14).
@@ 2.2 Mathematical Formalization
A trajectory is the complete sequence $\tau=(S_{0},A_{0},S_{1},\ldots,A_{H-1},S_{H})$. We use uppercase letters for random variables; realized samples are denoted $\tau^{(i)}=(s_{0}^{(i)},a_{0}^{(i)},\ldots,s_{H}^{(i)})$.
@@ 2.3 Supervision Models
The learner sees what happened at each step but not whether each step was correct.
@@
$\mathcal{D}_{\mathrm{proc}}=\{(\tau^{(i)},r_{0}^{(i)},\ldots,r_{H-1}^{(i)},R^{(i)})\}_{i=1}^{n},$ (3)
@@
$\mathcal{D}_{k}=\{(\tau^{(i)},r_{k}^{(i)},r_{2k}^{(i)},\ldots,r_{\lfloor(H-1)/k\rfloor\cdot k}^{(i)},R^{(i)})\}_{i=1}^{n},$ (4)
where the labeled steps are $\{t\in\{1,\ldots,H-1\}:t\equiv 0\pmod{k}\}$. This interpolates between near-process supervision ($k=1$, all steps except $r_{0}$ labeled) and outcome supervision ($k\geq H$, no step labels). Omitting $r_{0}$ is without loss for our results; it ensures that coarse supervision provides no step-level information.
@@ 2.4 Information-Theoretic Preliminaries
$\chi^{2}(P\|Q):=\int\frac{(dP-dQ)^{2}}{dQ}=\mathbb{E}_{Q}\left[\left(\frac{dP}{dQ}-1\right)^{2}\right].$ (5)
@@
$\eta_{\chi^{2}}(K):=\sup_{\begin{subarray}{c}P,Q:P\ll Q\\ 0<\chi^{2}(P\|Q)<\infty\end{subarray}}\frac{\chi^{2}(PK\|QK)}{\chi^{2}(P\|Q)},$ (6)
@@
$\chi^{2}(P_{S_{H}}\|Q_{S_{H}})\leq\left(\prod_{t=0}^{H-1}\eta_{\chi^{2}}(K_{t})\right)\chi^{2}(P_{S_{0}}\|Q_{S_{0}}).$ (8)
@@ 3 The Signal Decay Bound: Why Early Steps Are Hardest
We now develop our first main result: the Signal Decay Bound. In settings ranging from manufacturing quality control to service analytics, a decision-maker observes process outcomes and must infer which intermediate stages were responsible for success or failure. The central question is: how much signal about an early step survives propagation through the remaining stages? We prove that this signal decays exponentially with distance from the step to the outcome, making credit assignment to early steps fundamentally harder than to late ones.
The section proceeds in three steps. Section 3.1 introduces the state abstraction that maps observable states to a coarser space on which the dynamics are Markov. Section 3.2 shows that the many-to-one structure of this abstraction induces contraction in the abstract transition kernels, quantifies the contraction coefficient, and states the standing assumption for subsequent results. Section 3.3 then proves the Signal Decay Bound, derives the critical horizon, and establishes tightness. Proofs of all results in this section appear in Appendix 10.
@@ 3.1 State Abstraction and Abstract Dynamics
$\mathcal{L}(\phi(S_{t+1})\mid S_{t}=s)=\mathcal{L}(\phi(S_{t+1})\mid S_{t}=s^{\prime}).$ (10)
@@
Lumpability is a modeling choice rather than a restrictive hypothesis: one defines $\phi$ so that the abstract dynamics are Markov. If an initial choice of $\phi$ fails the condition, one refines the partition by augmenting $\phi$ with additional context until the condition holds, at the cost of a larger abstract space. The interesting question is not whether lumpability holds (it can always be arranged) but how much the resulting abstract kernels $K_{t}$ mix: do different abstract states lead to overlapping or distinct distributions over successor states? A coarser abstraction lumps more observable states together, increasing overlap and accelerating information loss. A finer abstraction preserves more distinctions but may require a larger state space. Section 3.2 formalizes this tradeoff. Section 14 in the Electronic Companion develops a concrete instantiation for language model reasoning.
@@ 3.2 Why Abstract Dynamics Contract
The following two propositions formalize the relationship between row overlap and contraction, providing concrete sufficient conditions for $\eta<1$.
@@
Consider a transition kernel $K$ on an abstract state space with $m$ states. If $K$ spreads probability such that $K(z^{\prime}\mid z)\geq p_{\min}>0$ for all state pairs $(z,z^{\prime})$, then $\eta_{\chi^{2}}(K)\leq 1-m\cdot p_{\min}$.
Proposition 3.4 provides a sufficient but strong condition: requiring $K(z^{\prime}\mid z)\geq p_{\min}$ for all state pairs is often unrealistic. The following proposition gives a weaker condition based on the Dobrushin coefficient, requiring only that each pair of rows has some overlap.
Proposition 3.5 (Contraction from Row Overlap)
@@
The resulting bound on $\eta$ is looser than Proposition 3.4 but applies more broadly, since it requires only pairwise row overlap rather than full minorization. In practice, $\alpha(K)$ can be estimated by sampling continuations from different prefixes and measuring their distributional overlap. In manufacturing, the same bound is computed directly from shop-floor data: bin each stage’s output into grades via $\phi$, estimate $K$ from empirical adjacent-stage transition frequencies, and compute $\alpha(K)$ from the resulting matrix.
Example 3.6 (Manufacturing Grade Transitions)
Consider a three-grade abstraction $Z_{t}\in\{\mathrm{A},\mathrm{B},\mathrm{C}\}$ where $\mathrm{A}$ denotes “within specification,” $\mathrm{B}$ denotes “marginal/reworkable,” and $\mathrm{C}$ denotes “defective,” in a setting with rework and defect masking. A downstream step such as polishing followed by retest can partially repair marginal units while occasionally degrading good ones. A representative grade-to-grade kernel is
$K=\begin{pmatrix}0.85&0.14&0.01\\ 0.55&0.35&0.10\\ 0.20&0.30&0.50\end{pmatrix}$,
where rows index the incoming grade and columns the outgoing grade. The minimum pairwise row overlap is achieved by the $\mathrm{A}$ and $\mathrm{C}$ rows:
[…]
so $\alpha(K)=0.35$ and Proposition 3.5 gives $\eta_{\chi^{2}}(K)\leq 1-\alpha(K)=0.65$. Even when the underlying physical transformation is repeatable, coarse grading and partial defect masking produce strict per-stage contraction on the abstract state space.
The preceding examples and propositions show that contraction ($\eta<1$) arises whenever the rows of the abstract kernel have nontrivial overlap. The only abstract chains that avoid contraction entirely are those whose transition matrix is a permutation, so that every state has a unique successor. We now state this as a standing assumption.
@@
Propositions 3.4 and 3.5 provide concrete upper bounds on $\eta$ in terms of the abstract kernel’s row overlap; Section 14 in the Electronic Companion shows how sampling temperature controls $\eta$ in language model reasoning.
Under Assumption 3.6, consider two hypotheses $H_{0}$ and $H_{1}$ that induce different distributions $P^{(0)}_{Z_{t}}$ and $P^{(1)}_{Z_{t}}$ over the abstract state at step $t$. Information about step $t$ must propagate through HtH-t transitions to reach the outcome:
$Z_{t}\xrightarrow{K_{t}}Z_{t+1}\xrightarrow{K_{t+1}}\cdots\xrightarrow{K_{H-1}}Z_{H}\to R.$
At each transition, the SDPI contracts the divergence by at most factor $\eta$, yielding:
$\text{At step }t:$ $\quad\chi^{2}(P^{(0)}_{Z_{t}}\|P^{(1)}_{Z_{t}})=\Delta^{2}$
$\text{At step }t+1:$ $\quad\chi^{2}(P^{(0)}_{Z_{t+1}}\|P^{(1)}_{Z_{t+1}})\leq\eta\cdot\Delta^{2}$
$\;\;\vdots$
$\text{At step }H:$ $\quad\chi^{2}(P^{(0)}_{Z_{H}}\|P^{(1)}_{Z_{H}})\leq\eta^{H-t}\cdot\Delta^{2}.$
After HtH-t transitions, the divergence shrinks by factor $\eta^{H-t}$. When $\eta<1$, this decay is exponential: the outcome distributions under $H_{0}$ and $H_{1}$ become nearly identical, requiring exponentially many samples to distinguish which hypothesis is true. The next subsection makes this precise.

### v2

@@ 1 Introduction
The challenge of attributing outcomes to early stages under terminal-only observation has gained particular urgency in AI, where large language models generate extended reasoning chains but only the final answer can be verified. Empirically, reinforcement learning (RL) value networks trained on outcome rewards “collapse”: they converge to near-constant predictions for early steps, as if outcomes carry no information about the chain’s beginning (Kazemnejad et al. 2024). Process reward models that provide step-level supervision dramatically improve performance (Lightman et al. 2024), but require expensive human annotation. When is such annotation necessary, and when does outcome supervision suffice?
This paper answers these questions by establishing information-theoretic foundations for credit assignment. Using Strong Data Processing Inequalities (SDPIs), we prove that signals connecting early steps to terminal outcomes decay exponentially with depth. The analysis reveals a sharp threshold: there exists a critical horizon $H_{\mathrm{crit}}$ such that steps within this distance of the outcome can be evaluated from endpoint data, while steps beyond require a number of samples that grows exponentially in the gap HtH-t. This is not an algorithmic limitation but an information-theoretic barrier. The critical horizon scales logarithmically with sample size and signal strength but inversely with a per-stage contraction rate that captures how much each processing step blurs the distinction between good and bad earlier steps. In AI reasoning systems, sampling temperature controls this blurring: higher temperature produces more diverse continuations, making it harder to trace outcomes back to early steps. In manufacturing, the analogous driver is the overlap among downstream quality-grade distributions: variability-reducing steps (polishing, annealing, rework), pooling operations (lot mixing, kitting, assembly), and coarse metrology bins all cause products of different upstream grades to produce similar distributions of grades at the next stage, fading the signal about the original grade. The critical horizon formula thus determines whether intermediate feedback is a cost-saving measure or a feasibility requirement.
@@
This many-to-one mapping from observations to abstract states is the mechanism behind the per-stage contraction introduced above: when multiple observable states collapse to the same abstract state, transitions from distinct starting points lead to overlapping distributions over next states, and statistical distinguishability shrinks. This occurs even when raw dynamics are deterministic (as with token generation in language models), because the induced dynamics on the abstract space contract. The implication for learning systems is immediate: when signal decay is severe, outcome distributions conditioned on different early states become nearly identical, and training data contains almost no information about which early states are better. Under the assumptions of our model, the critic collapse phenomenon observed in AI systems (Kazemnejad et al. 2024) is statistically optimal behavior given the available information, and not a bug to be fixed but a symptom of a fundamental limit.
A natural response to signal decay is to generate more observations: multiple parallel trajectories from the same starting point, as in replicated testing or width-based methods like GRPO (Group Relative Policy Optimization; DeepSeek-AI 2024). We show this strategy provides only limited relief: correlation among rollouts caps the effective width at $1/\rho$, where $\rho$ is the pairwise correlation between rollout outcomes, and adding more rollouts yields diminishing returns.
@@
We develop a detailed application to large language model reasoning (Section 18 in the Electronic Companion), where the state abstraction maps token sequences to semantic content and sampling temperature controls the contraction coefficient. We validate the theoretical predictions through numerical experiments on both synthetic Markov chains and LLM reasoning chains on GSM8K (a grade-school math benchmark); full experimental details appear in Appendix 17.
@@
The framework yields concrete guidance for both AI and operations practitioners. For AI systems, it explains why process reward models improve performance (they provide signal beyond the critical horizon), why methods like VinePPO (value-based intermediate-step process supervision; Kazemnejad et al. 2024) succeed (they reduce effective depth below $H_{\mathrm{crit}}$), and why width-based methods face diminishing returns (correlation caps effective width). For operations managers, it provides a direct diagnostic. Consider a 50-stage semiconductor process. With contraction rate $\eta=0.85$, signal strength $\Delta^{2}=0.2$, and 10,000 test wafers, $H_{\mathrm{crit}}\approx 47$ (simplified; $\approx 48$ with $\epsilon=0.1$): at least one intermediate checkpoint is required. With $\eta=0.90$, the critical horizon extends to 72 steps, and terminal testing alone suffices. (In practice, $\eta$ is auditable: estimate adjacent-stage grade transition matrices from production data and compute the overlap-based bounds in Section 3.2.)
@@ 1.1 Literature Review
Our results sit at the intersection of inspection design in operations research, outcome-versus-process supervision in deep sequential decision systems, and information-theoretic limits for Markovian signal propagation. On the OR side, we build on classical checkpoint placement models (Lindsay and Bishop 1964, Raz 1986, Tapiero 1987), but focus on a different binding constraint: observability under terminal-only evaluation in deep processes, rather than expected-cost tradeoffs under persistent defect signals. On the AI side, recent work documents that outcome-only supervision can fail on long reasoning chains and that step-level supervision can help (Lightman et al. 2024, Uesato et al. 2022, Kazemnejad et al. 2024); we provide a non-asymptotic, information-theoretic characterization of when outcome supervision becomes insufficient. In standard RL settings with bounded concentrability, outcome and process supervision are polynomially equivalent (Jia et al. 2025); our setting falls outside this regime because concentrability grows exponentially in horizon. Independent concurrent work by chen2025exponential establishes an exponential separation using function approximation complexity, complementing our information-theoretic approach.
More broadly, this work contributes to a growing OR-for-AI agenda that applies operations research methodology to AI system design (Snell et al. 2025, Besta et al. 2025, Elmachtoub and Grigas 2022, Wiberg et al. 2025), here showing that classical tools such as strong data processing inequalities and minimax inspection design yield precise answers to questions the AI community has addressed primarily through empirical heuristics. Technically, the analysis leverages strong data processing inequalities (Polyanskiy and Wu 2017, Raginsky 2016) and Le Cam-style hypothesis testing lower bounds (Le Cam 1973) to establish tight sample-complexity limits and the closed-form critical horizon. A detailed, stream-by-stream literature discussion (inspection design, RL theory, information theory, and OR-for-AI) is provided in the Electronic Companion (Appendix 10).
@@ 1.2 Paper Organization
Section 6 synthesizes these results into optimal inspection schedules: uniform spacing under homogeneous contraction and a greedy information-distance algorithm under heterogeneous contraction. Section 7 provides a concise summary of numerical validation; full experimental setups and all plots are deferred to the Electronic Companion (Appendix 17). The Electronic Companion also develops a detailed application to language model reasoning (Section 18).
@@ 2.2 Mathematical Formalization
A trajectory is the complete sequence $\tau=(S_{0},A_{0},S_{1},\ldots,A_{H-1},S_{H})$. We use uppercase letters for random variables; realized samples are denoted $\tau^{(i)}=(s_{0}^{(i)},a_{0}^{(i)},\ldots,s_{H}^{(i)})$. In sequential validity tasks, terminal success requires every step to be correct: $R=\prod_{t=0}^{H-1}r_{t}$; this structural relationship motivates the multiplicative objective introduced in Section 5.
@@ 2.3 Supervision Models
The learner sees what happened at each step but not whether each step was correct. Consequently, the only supervision signal about step validity is through $R$, and any test for whether step $t$ was correct must rely on the induced separation in outcome distributions; this is why the lower bounds in Section 3 are stated in terms of $P^{(0)}_{R}$ versus $P^{(1)}_{R}$.
@@
$\mathcal{D}_{\mathrm{proc}}=\{(\tau^{(i)},r_{0}^{(i)},\ldots,r_{H-1}^{(i)},R^{(i)})\}_{i=1}^{n},$}
@@
$\mathcal{D}_{k}=\{(\tau^{(i)},r_{k}^{(i)},r_{2k}^{(i)},\ldots,r_{\lfloor(H-1)/k\rfloor\cdot k}^{(i)},R^{(i)})\}_{i=1}^{n},$}
That is, labels are provided at steps $t=k,2k,\ldots$ (we omit $t=0$ by convention). This interpolates between near-process supervision ($k=1$, all steps except $r_{0}$ labeled) and outcome supervision ($k\geq H$, no step labels). The convention ensures that coarse supervision provides no step-level information.
@@ 2.4 Information-Theoretic Preliminaries
$\chi^{2}(P\|Q):=\int\frac{(dP-dQ)^{2}}{dQ}=\mathbb{E}_{Q}\left[\left(\frac{dP}{dQ}-1\right)^{2}\right].$}
@@
$\eta_{\chi^{2}}(K):=\sup_{\begin{subarray}{c}P,Q:P\ll Q\\ 0<\chi^{2}(P\|Q)<\infty\end{subarray}}\frac{\chi^{2}(PK\|QK)}{\chi^{2}(P\|Q)},$}
@@
$\chi^{2}(P_{S_{H}}\|Q_{S_{H}})\leq\left(\prod_{t=0}^{H-1}\eta_{\chi^{2}}(K_{t})\right)\chi^{2}(P_{S_{0}}\|Q_{S_{0}}).$}
@@ 3 The Signal Decay Bound: Why Early Steps Are Hardest
We now prove that information about step $t$ decays exponentially as it propagates through the remaining HtH-t stages, making credit assignment to early steps fundamentally harder than to late ones. We first introduce a state abstraction under which abstract dynamics are Markov (Section 3.1), show that the many-to-one structure of this abstraction induces contraction (Section 3.2), and then prove the Signal Decay Bound, derive the critical horizon, and establish tightness (Section 3.3). Proofs appear in Appendix 11.
@@ 3.1 State Abstraction and Abstract Dynamics
$\mathcal{L}(\phi(S_{t+1})\mid S_{t}=s)=\mathcal{L}(\phi(S_{t+1})\mid S_{t}=s^{\prime}).$}
@@
Lumpability is a modeling choice rather than a restrictive hypothesis: one defines $\phi$ so that the abstract dynamics are Markov. If an initial choice of $\phi$ fails the condition, one refines the partition by augmenting $\phi$ with additional context until the condition holds, at the cost of a larger abstract space. The interesting question is not whether lumpability holds (it can always be arranged, at the cost of enlarging the abstract state space) but how much the resulting abstract kernels $K_{t}$ mix: do different abstract states lead to overlapping or distinct distributions over successor states? A coarser abstraction lumps more observable states together, increasing overlap and accelerating information loss. A finer abstraction preserves more distinctions but may require a larger state space. Section 3.2 formalizes this tradeoff. Section 18 in the Electronic Companion develops a concrete instantiation for language model reasoning.
@@ 3.2 Why Abstract Dynamics Contract
The following two propositions formalize the relationship between row overlap and $\chi^{2}$-contraction, providing concrete sufficient conditions for $\eta<1$. Both build on classical Markov chain mixing techniques (Doeblin decomposition, Dobrushin’s ergodicity coefficient) combined with the SDPI framework (Lemma 2.9); proofs appear in Appendix 11.
@@
Consider a transition kernel $K$ on an abstract state space with $|\mathcal{Z}|$ states. If $K$ spreads probability such that $K(z^{\prime}\mid z)\geq p_{\min}>0$ for all state pairs $(z,z^{\prime})$, then $\eta_{\chi^{2}}(K)\leq 1-|\mathcal{Z}|\cdot p_{\min}$ (note $|\mathcal{Z}|\cdot p_{\min}\leq 1$ since rows sum to one).
Proposition 3.4 provides a sufficient but strong condition: requiring $K(z^{\prime}\mid z)\geq p_{\min}$ for all state pairs is often unrealistic. The following proposition gives a weaker condition based on the Dobrushin coefficient (dobrushin1956central), requiring only that each pair of rows has some overlap.
Proposition 3.5 (Contraction from Row Overlap (cf. dobrushin1956central, Polyanskiy and Wu 2017))
@@
The resulting bound on $\eta$ is looser than Proposition 3.4 but applies more broadly, since it requires only pairwise row overlap rather than full minorization. In practice, $\alpha(K)$ can be estimated by sampling continuations from different prefixes and measuring their distributional overlap. In manufacturing, the same bound is computed from existing instrumentation or pilot audits: bin each stage’s output into grades via $\phi$, estimate $K$ from empirical adjacent-stage transition frequencies, and compute $\alpha(K)$ from the resulting matrix, giving $\eta\leq 1-\alpha(K)$ (Proposition 3.5). Example 15.1 in Appendix 15 illustrates this computation for a three-grade quality transition kernel.
Example 3.3, Example 15.1 in Appendix 15, and Propositions 3.43.5 show that contraction ($\eta<1$) arises whenever the rows of the abstract kernel have nontrivial overlap. Contraction fails (i.e., $\eta=1$ remains possible) only when some pairs of rows have disjoint support ($\alpha(K)=0$), as in deterministic dynamics with distinct successors. We now formalize the assumption that contraction is strict.
@@
Propositions 3.4 and 3.5 provide concrete upper bounds on $\eta$ in terms of the abstract kernel’s row overlap. In Section 18 of the Electronic Companion, we establish how sampling temperature governs $\eta$ in language model reasoning, connecting the abstract contraction framework to a concrete and practically important design lever.